1. Briefly describe the concepts of ROI, Payback Period, and Breakeven Analysis

2. Company XYZ is looking to launch a new product next year, Product A will cost $50,000 to launch, and will bring in a profit of $100,000, Product B will cost $500,000 to launch and bring in a profit of $900,000. What is the ROI of each product and which should the company launch?

3. Project A which yields the following cash flows over its five year life. What is the Payback?

Year Cash Flow Net Cash Flow
0 -1,000 -1,000
1 500 -500
2 400 -100
3 200 100
4 200 300
5 100 400

4. A firm produces and sells a product with these values:

F = Total Fixed costs = $1,200 V = Variable cost per unit = $15P = Selling price per unit = $40

What is the breakeven point in Units? How many Units does it need to sell in order to make $1,300?
Answer #1

Parrt-2:

Particulars

Product-A

Prodcut-B

Cost

           50,000

       5,00,000

Profit

       1,00,000

       9,00,000

ROI (Profit/Cost)*100

200%

180%

Based on ROI, XYZ should launch Product-A

.

Part-3:

Year

Cash Flows

Cummulative Cash Flows

0

-1000

-1000

1

500

-500

2

400

-100

3

200

100

(2+100/200)

4

200

300

5

100

400

.

Payback of Project-A = 2 + (100/200)

= 2.5 years

.

Part-4:

Particulars

Amount ($)

Seeling Price p.u.

                    40

Less: Variable Cost p.u.

                    15

Contribution p.u.

                    25

Fixed Cost

              1,200

Break Even Point

                    48

(1200/25)

Units to sell in order to make $1,300 profit = (Profit + Fixed Cost)/Contribution p.u

= (1300+1200)/25

= 2500/25

= 100 units